### v3

# A generic transformation is invertible

Abstract.
We show that, on a standard non-atomic probability space, invertible measure-preserving transformations form a dense $G_{\delta}$ subset of the space of all measure-preserving transformations endowed with the strong (==weak) operator topology. This implies that all properties which are generic for invertible transformations are also generic for general ones.
Key words and phrases:
measure-preserving transformations, generic, invertible
2010 Mathematics Subject Classification:
28D05, 37A05, 54H05
1. Introduction
Let $(X,\mu)$ be a standard non-atomic probability space and $T:X\to X$ be measure preserving, i.e., satisfies $\mu(T^{-1}(A))=\mu(A)$ for every measurable $A\subset X$. Recall that the corresponding Koopman operator [^1] We denote it by the same letter as the transformation following a tradition in ergodic theory. $T$ on $L^{2}(X,\mu)$ is defined by $(Tf)(x):=f(Tx)$ and is an isometry on $L^{2}(X,\mu)$. Moreover, this operator is invertible, i.e., unitary, if and only if222see, e.g., [7, Rem. 8.48(d)](e) for the “only if” direction the underlying transformation is essentially invertible.
Denote by $\mathcal{T}$ the set of all measure-preserving transformations on $(X,\mu)$. Then $\mathcal{T}$ is a complete metric [^3] An example of a metric inducing this topology is $d(T,S):=\sum_{j=1}^{\infty}\frac{\|Tf_{j}-Sf_{j}\|}{2^{j}\|f_{j}\|}$ for a fixed dense subset $(f_{j})$ of $L^{2}(X,\mu)\setminus\{0\}$. space with respect to the strong (==weak) operator topology for the corresponding Koopman operators, cf. Halmos [12].
Usually one works with the space $\mathcal{T}_{\text{inv}}$ of all invertible measure-preserving transformations on $(X,\mu)$ with respect to the strong (==weak) operator topology [^4] One says that a net $(T_{\alpha})$ of operators on a Hilbert space converges strongly to an operator $T$ if $\lim_{\alpha}T_{\alpha}=T$ and $\lim_{\alpha}T_{\alpha}^{*}=T^{*}$ strongly. for the corresponding Koopman operators which is also a complete metric space, see Halmos [12], and studies generic properties of transformations therein. Since the existence proof by Halmos [11] and Rokhlin [18] of a weakly mixing but not mixing invertible transformation (with a first concrete example constructed much later by Chacon [2]), the study of generic properties in $\mathcal{T}_{\text{inv}}$ has been an active and exciting area of research. For example, the following further properties are known to be generic in $\mathcal{T}_{\text{inv}}$: zero entropy (Rohlin [19]), rigidity (Katok [13], Nadkarni [17, Chapter 8]), having roots of all orders (see King [15]) and embedding into many flows (see de la Rue, de Sam Lazaro [4], Stepin, Eremenko [29]). For more results see, e.g., Stepin [28], Choksi, Nadkarni [3], Nadkarni [17], del Junco, Lemańczyk [5], Ageev [1], Ryzhikov [21, 24], Solecki [26, 27]. For generic properties of extensions we refer to Schnurr [25], Glasner, Weiss [8], Ryzhikov [22, 23], Kozyreva, Ryzhikov [16].
Non-invertible transformations are harder to work with than invertible ones and are not so well-studied. We will show that nevertheless all generic properties from $\mathcal{T}_{\text{inv}}$ transfer to $\mathcal{T}$ by the following.
Theorem 1.1 (Invertibility is generic).
The set $\mathcal{T}_{\text{inv}}$ of all invertible measure-preserving transformations is a dense $G_{\delta}$ subset of $\mathcal{T}$.
In particular, the set $\mathcal{T}_{\text{inv}}$ is residual in $\mathcal{T}$. After the proof of the dense $G_{\delta}$ property in Section 2 we also present an alternative, more elementary and direct proof of residuality in Section 3.
An operator-theoretic analogue of Theorem 1.1 with a residuality statement instead of the dense $G_{\delta}$ property was shown in [6], with the corresponding $G_{\delta}$-refinement essentially done by Grivaux, Matheron, Menet [10].
Acknowledgement. The author is indebted to Valentin Gillet for very helpful discussions regarding the points of continuity method in the operator theoretic context and, in particular, for the references [9, 10].
2. Proof
A linear operator on $L^{2}(X,\mu)$ is called bi-stochastic if $P\mathbf{1}=P^{*}\mathbf{1}=\mathbf{1}$ and $P$ is positive, i.e., satisfies $Pf\geq 0$ for every $f\geq 0$. Recall that such operators are contractive on $L^{\infty}(X,\mu)$, $L^{1}(X,\mu)$ and hence by interpolation on every $L^{p}(X,\mu)$, $p\in[1,\infty]$.
The following result describes the weak closure of $\mathcal{T}_{\text{inv}}$, see Vershik [30, Sections 1–2] and a simplification in Ryzhikov [20, Remark after Lemma 5.5]. We present the details for the reader’s convenience.
Lemma 2.1.
Invertible measure-preserving transformations are dense in the set of all bi-stochastic operators on $L^{2}(X,\mu)$ with respect to the weak operator topology.
Proof.
Assume without loss of generality that $X=[0,1]$ and $\mu$ is the Lebesgue measure, see, e.g., [7, Rem. 8.48(d)]. Let $P$ be a bi-stochastic operator on $L^{2}[0,1]$, $n\in\mathbb{N}$ and divide $[0,1]$ into $2^{n}$ dyadic intervals $I_{j}:=\left[\frac{j-1}{2^{n}},\frac{j}{2^{n}}\right]$, $j\in\{1,\ldots,2^{n}\}$. We call the characteristic functions of such intervals dyadic functions of degree $n$. Denote $a_{jk}:=\langle P\mathbf{1}_{I_{j}},\mathbf{1}_{I_{k}}\rangle$ for $j,k\in\{1,\ldots,2^{n}\}$. Since $P$ is bi-stochastic, the matrix $(2^{n}a_{jk})$ is bi-stochastic, i.e., $a_{jk}\geq 0$ and $\sum_{j}a_{jk}=\sum_{k}a_{jk}=\frac{1}{2^{n}}$ holds for all $j,k$. It is now easy to construct a measure-preserving transformation $T$ on $L^{2}[0,1]$ satisfying
(1) $\langle T\mathbf{1}_{I_{j}},\mathbf{1}_{I_{k}}\rangle=\langle P\mathbf{1}_{I_{j}},\mathbf{1}_{I_{k}}\rangle\quad\forall j,k\in\{1,\ldots,2^{n}\}$
as follows. Divide the interval $I_{1}$ into subintervals $I_{1k}$ of length $a_{1k}$, $k\in\{1,\ldots,2^{n}\}$, which we count from left to right. Define $T$ so that it leaves the first subinterval $I_{11}$ as it is, shifts the second subinterval $I_{12}$ to the beginning of $I_{2}$, etc. Analogously define $T$ on $I_{2}$ shifting the subinterval $I_{21}$ of length $a_{21}$ to the beginning of the not yet covered part of $I_{1}$ (intersecting the covered part in one end point), shifting the subinterval of length $a_{22}$ to the left of the not yet covered part of $I_{2}$ etc. In this way, for every $n$ we obtain an invertible measure-preserving transformation $T_{n}$ on $L^{2}[0,1]$ satisfying (1). Note that by (1) and linearity $\langle Tf,g\rangle=\langle Pf,g\rangle$ holds for all $f,g$ being linear combinations of dyadic functions of degree $\leq n$. Since the linear span of dyadic functions is dense in $L^{2}[0,1]$, the sequence $(T_{n})$ approximates $P$ weakly. ∎
Remark.
We will not need the full strength of Lemma 2.1 but just the fact that invertible transformations are dense in $\mathcal{T}$.
The following is an ergodic theoretic analogue of Grivaux, Matheron, Menet [10, Prop. 2.11]. Here and later we denote by SOT and $\texttt{SOT}^{*}$ the strong and the strong operator topology on the Koopman operators, respectively.
Lemma 2.2.
The set $\mathcal{T}_{\text{inv}}$ coincides with the set of continuity points of the identity map $\text{id}:(\mathcal{T},\texttt{SOT})\to(\mathcal{T},\texttt{SOT}^{*})$.
Proof.
Let $T\in\mathcal{T}_{\text{inv}}$ and $(T_{n})\subset\mathcal{T}$ converging to $T$ in the strong operator topology. Then it converges also in the weak operator topology implying $T_{n}^{*}\to T^{*}$ weakly. Since $T$ is by assumption invertible, $T^{*}$ is an isometry as well and hence [^5] Recall that weak (operator) convergence of contractions on a Hilbert space implies strong convergence as soon as the limit operator is an isometry. $T_{n}^{*}\to T^{*}$ strongly. Thus $T$ is a point of continuity of $\text{id}:(\mathcal{T},\texttt{SOT})\to(\mathcal{T},\texttt{SOT}^{*})$.
Let now $T\in\mathcal{T}$ be such a point of continuity and let $f\in L^{2}(X,\mu)$. By Lemma 2.1 there exists a sequence $(T_{n})\subset\mathcal{T}_{\text{inv}}$ converging to $T$ weakly and hence strongly. By the point of continuity assumption we have
$\|f\|=\lim_{n\to\infty}\|T_{n}^{*}f\|=\|T^{*}f\|,$
and in particular $T^{*}$ is injective. Equivalently, $T$ has dense range, so [^6] Recall that isometries always have closed range. is surjective and hence invertible. This shows $T\in\mathcal{T}_{\text{inv}}$. ∎
The idea of the following proof in the operator theoretic setting, being a simplification of the one by Grivaux, Matheron, Menet [9, beginning of Sect. 3], was communicated to the author by Valentin Gillet.
Lemma 2.3.
The space $(\mathcal{T},\texttt{SOT}^{*})$ is separable.
Proof.
Let $(f_{n})_{n=1}^{\infty}$ be dense in $L^{2}(X,\mu)$ and consider the set
$A:=\{((Tf_{k})_{k=1}^{\infty},(T^{*}f_{k})_{k=1}^{\infty}):\,T\in\mathcal{T}\}\subset(L^{2}(X,\mu))^{\mathbb{N}}\times(L^{2}(X,\mu))^{\mathbb{N}}=:Y.$
Observe first that $Y$ is a Polish space, where $(L^{2}(X,\mu))^{\mathbb{N}}$ is endowed with the product topology induced by the norm topology on $L^{2}(X,\mu)$. Since for metric spaces separability is equivalent to second countability, $A$ is separable as well.
Thus let $(a_{n})$ be dense in $A$ and denote by $(T_{n})$ the corresponding countable subset of $\mathcal{T}$. We show that $(T_{n})$ is dense in $(\mathcal{T},\texttt{SOT}^{*})$. Let $T\in\mathcal{T}$ and let $(n_{l})$ be a subsequence of $\mathbb{N}$ such that $(a_{n_{l}})$ approximates $((Tf_{k})_{k=1}^{\infty},(T^{*}f_{k})_{k=1}^{\infty})$ in $Y$. This means that $(T_{n_{l}})$ approximates $T$ and $(T^{*}_{n_{l}})$ approximates $T^{*}$ in SOT, i.e., $(T_{n_{l}})$ approximates $T$ in $\texttt{SOT}^{*}$. The proof is complete. ∎
Proof of Theorem 1.1.
By a classical result of Baire, see Kechris [14, Thm. 24.14], the set of continuity points of a Baire class $1$ map [^7] A map $f:X\to Y$ between topological spaces $X$ and $Y$ is called of Baire class $1$ (or Borel $1$) if the preimage of every open subset of $Y$ is $F_{\sigma}$ in $X$. $f:X\to Y$ form a dense $G_{\delta}$ subset of $X$ whenever $X,Y$ are metrizable and $Y$ is separable. By Lemmata 2.2 and 2.3 it just remains to check the Baire class $1$ property of the identity map $\mathrm{id}:(\mathcal{T},\texttt{SOT})\to(\mathcal{T},\texttt{SOT}^{*})$. This can be done as in [10, Lemma 2.9] and we repeat here the simple argument for completeness.
Every $\texttt{SOT}^{*}$-open subset of $\mathcal{T}$ can be written as a countable union of finite intersections of sets either of the form
$U_{S,f,\varepsilon}:=\left\{T\in\mathcal{T}:\,\|(T-S)f\|<\varepsilon\right\}=\bigcup_{n=1}^{\infty}\left\{T\in\mathcal{T}:\,\|(T-S)f\|\leq\varepsilon-\frac{1}{n}\right\}$
or of the form
$U^{*}_{S,f,\varepsilon}:=\{T\in\mathcal{T}:\,\|(T^{*}-S^{*})f\|<\varepsilon\}=\bigcup_{n=1}^{\infty}\left\{T\in\mathcal{T}:\,\|(T^{*}-S^{*})f\|\leq\varepsilon-\frac{1}{n}\right\}$
for $S\in\mathcal{T}$, $f\in L^{2}(X,\mu)$ and $\varepsilon>0$. Since $\|(T^{*}-S^{*})f\|\leq\varepsilon-\frac{1}{n}$ can be written as
$|\langle(T-S)g,f\rangle|=|\langle g,(T^{*}-S^{*})f\rangle|\leq\varepsilon-\frac{1}{n}\qquad\forall g\in L^{2}(X,\mu)\text{ with }\|g\|=1,$
the SOT-$F_{\sigma}$ property of $\texttt{SOT}^{*}$-open sets is clear. Thus $\mathrm{id}:(\mathcal{T},\texttt{SOT})\to(\mathcal{T},\texttt{SOT}^{*})$ is of Baire class $1$. ∎
3. Alternative proof of residuality
We now present an alternative, more elementary and direct proof of the residuality of $\mathcal{T}_{\text{inv}}$ in $\mathcal{T}$ using Lemma 2.1 and argue as in [6]. Let $(f_{j})$ be dense in $L^{2}(X,\mu)$. Since every isometry has closed range,
(2) $\mathcal{T}\setminus\mathcal{T}_{\text{inv}}=\bigcup_{j,k\in\mathbb{N}}\left\{T\in\mathcal{T}:\,\text{dist}(f_{j},\text{rg}T)>\frac{1}{k}\right\}=:\bigcup_{j,k\in\mathbb{N}}M_{j,k}.$
We will show that each set $M_{j,k}$ is nowhere dense. By density of $\mathcal{T}_{\text{inv}}$ in $\mathcal{T}$ it is enough to show $\overline{M_{j,k}}\cap\mathcal{T}_{\text{inv}}=\emptyset$. Assume the opposite and let $T\in\mathcal{T}_{\text{inv}}$ satisfy $\lim_{n\to\infty}T_{n}=T$ strongly for some $(T_{n})\subset\mathcal{T}$ with $\text{dist}(f_{j},\text{rg}T_{n})>\frac{1}{k}$ for every $n\in\mathbb{N}$. Then $g:=T^{-1}f_{j}$ satisfies
$\lim_{n\to\infty}T_{n}g=Tg=f_{j},$
i.e., $\lim_{n\to\infty}\text{dist}(f_{j},\text{rg}T_{n})=0$, a contradiction.