### v4

# A generic transformation is invertible

Abstract.
We show that, on a standard non-atomic probability space, invertible measure-preserving transformations form a dense $G_{\delta}$ subset of the space of all measure-preserving transformations endowed with the strong (==weak) operator topology. This implies that all properties which are generic for invertible transformations are also generic for general ones.
Key words and phrases:
measure-preserving transformations, generic, invertible
2010 Mathematics Subject Classification:
28D05, 37A05, 54H05
1. Introduction
Let $(X,\mu)$ be a standard non-atomic probability space and $T:X\to X$ be measure preserving, i.e., satisfies $\mu(T^{-1}(A))=\mu(A)$ for every measurable $A\subset X$. Recall that the corresponding Koopman operator [^1] We denote it by the same letter as the transformation following a tradition in ergodic theory. $T$ on $L^{2}(X,\mu)$ is defined by $(Tf)(x):=f(Tx)$ and is an isometry on $L^{2}(X,\mu)$. Moreover, this operator is invertible, i.e., unitary, if and only if222see, e.g., [7, Rem. 8.48(e)] for the “only if” direction the underlying transformation is essentially invertible.
Denote by $\mathcal{T}$ the set of all measure-preserving transformations on $(X,\mu)$. Then $\mathcal{T}$ is a complete metric [^3] An example of a metric inducing this topology is $d(T,S):=\sum_{j=1}^{\infty}\frac{\|Tf_{j}-Sf_{j}\|}{2^{j}\|f_{j}\|}$ for a fixed dense subset $(f_{j})$ of $L^{2}(X,\mu)\setminus\{0\}$. space with respect to the strong (==weak) operator topology for the corresponding Koopman operators, cf. Halmos [12].
Usually one works with the space $\mathcal{T}_{\text{inv}}$ of all invertible measure-preserving transformations on $(X,\mu)$ with respect to the strong (==weak) operator topology [^4] One says that a net $(T_{\alpha})$ of operators on a Hilbert space converges strongly to an operator $T$ if $\lim_{\alpha}T_{\alpha}=T$ and $\lim_{\alpha}T_{\alpha}^{*}=T^{*}$ strongly. for the corresponding Koopman operators which is also a complete metric space, see Halmos [12], and studies generic properties of transformations therein. Since the existence proof by Halmos [11] and Rokhlin [18] of a weakly mixing but not mixing invertible transformation (with a first concrete example constructed much later by Chacon [2]), the study of generic properties in $\mathcal{T}_{\text{inv}}$ has been an active and exciting area of research. For example, the following further properties are known to be generic in $\mathcal{T}_{\text{inv}}$: zero entropy (Rohlin [19]), rigidity (Katok [13], Nadkarni [17, Chapter 8]), having roots of all orders (see King [15]) and embedding into many flows (see de la Rue, de Sam Lazaro [4], Stepin, Eremenko [29]). For more results see, e.g., Stepin [28], Choksi, Nadkarni [3], Nadkarni [17], del Junco, Lemańczyk [5], Ageev [1], Ryzhikov [21, 24], Solecki [26, 27]. For generic properties of extensions we refer to Schnurr [25], Glasner, Weiss [8], Ryzhikov [22, 23], Kozyreva, Ryzhikov [16].
Non-invertible transformations are harder to work with than invertible ones and are not so well-studied. We will show that nevertheless all generic properties from $\mathcal{T}_{\text{inv}}$ transfer to $\mathcal{T}$ by the following.
Theorem 1.1 (Invertibility is generic).
The set $\mathcal{T}_{\text{inv}}$ of all invertible measure-preserving transformations is a dense $G_{\delta}$ subset of $\mathcal{T}$.
In particular, the set $\mathcal{T}_{\text{inv}}$ is residual in $\mathcal{T}$. We present two proofs of the $G_{\delta}$-property, a short one in Section 3 based on the fact that an isometry is invertible if and only if it has dense range, and a longer one in Section 4 using the points of continuity method.
An operator-theoretic analogue of Theorem 1.1 with a residuality statement instead of the dense $G_{\delta}$ property was shown in [6], with the corresponding $G_{\delta}$-refinement essentially done by Grivaux, Matheron, Menet [10].
Acknowledgements. The author is very grateful to Ethan Akin for a crystallisation of the argument in Section 3 from its original version concerning residuality. She is also indebted to Valentin Gillet for very helpful discussions regarding the points of continuity method in the operator theoretic context and, in particular, for the references [9, 10].
2. Proof: Density
A linear operator on $L^{2}(X,\mu)$ is called bi-stochastic if $P\mathbf{1}=P^{*}\mathbf{1}=\mathbf{1}$ and $P$ is positive, i.e., satisfies $Pf\geq 0$ for every $f\geq 0$. Recall that such operators are contractive on $L^{\infty}(X,\mu)$, $L^{1}(X,\mu)$ and hence by interpolation on every $L^{p}(X,\mu)$, $p\in[1,\infty]$.
The following result describes the weak closure of $\mathcal{T}_{\text{inv}}$, see Vershik [30, Sections 1–2] and a simplification in Ryzhikov [20, Remark after Lemma 5.5]. We present the details for the reader’s convenience.
Lemma 2.1.
Invertible measure-preserving transformations are dense in the set of all bi-stochastic operators on $L^{2}(X,\mu)$ with respect to the weak operator topology.
Proof.
Assume without loss of generality that $X=[0,1]$ and $\mu$ is the Lebesgue measure, see, e.g., [7, Rem. 8.48(d)]. Let $P$ be a bi-stochastic operator on $L^{2}[0,1]$, $n\in\mathbb{N}$ and divide $[0,1]$ into $2^{n}$ dyadic intervals $I_{j}:=\left[\frac{j-1}{2^{n}},\frac{j}{2^{n}}\right]$, $j\in\{1,\ldots,2^{n}\}$. We call the characteristic functions of such intervals dyadic functions of degree $n$. Denote $a_{jk}:=\langle P\mathbf{1}_{I_{j}},\mathbf{1}_{I_{k}}\rangle$ for $j,k\in\{1,\ldots,2^{n}\}$. Since $P$ is bi-stochastic, the matrix $(2^{n}a_{jk})$ is bi-stochastic, i.e., $a_{jk}\geq 0$ and $\sum_{j}a_{jk}=\sum_{k}a_{jk}=\frac{1}{2^{n}}$ holds for all $j,k$. It is now easy to construct a measure-preserving transformation $T$ on $L^{2}[0,1]$ satisfying
(1) $\langle T\mathbf{1}_{I_{j}},\mathbf{1}_{I_{k}}\rangle=\langle P\mathbf{1}_{I_{j}},\mathbf{1}_{I_{k}}\rangle\quad\forall j,k\in\{1,\ldots,2^{n}\}$
as follows. Divide the interval $I_{1}$ into subintervals $I_{1k}$ of length $a_{1k}$, $k\in\{1,\ldots,2^{n}\}$, which we count from left to right. Define $T$ so that it leaves the first subinterval $I_{11}$ as it is, shifts the second subinterval $I_{12}$ to the beginning of $I_{2}$, etc. Analogously define $T$ on $I_{2}$ shifting the subinterval $I_{21}$ of length $a_{21}$ to the beginning of the not yet covered part of $I_{1}$ (intersecting the covered part in one end point), shifting the subinterval of length $a_{22}$ to the left of the not yet covered part of $I_{2}$ etc. In this way, for every $n$ we obtain an invertible measure-preserving transformation $T_{n}$ on $L^{2}[0,1]$ satisfying (1). Note that by (1) and linearity $\langle Tf,g\rangle=\langle Pf,g\rangle$ holds for all $f,g$ being linear combinations of dyadic functions of degree $\leq n$. Since the linear span of dyadic functions is dense in $L^{2}[0,1]$, the sequence $(T_{n})$ approximates $P$ weakly. ∎
Remark.
We will not need the full strength of Lemma 2.1 but just the fact that invertible transformations are dense in $\mathcal{T}$.
3. First Proof of $G_{\delta}$ Property
The first proof of the $G_{\delta}$ property of $\mathcal{T}_{\text{inv}}$ in $\mathcal{T}$ is inspired by [6].
Let $(f_{j})$ be dense in $L^{2}(X,\mu)$. Since every isometry has closed range and $(Tf_{j})$ is dense in it, we have
$\mathcal{T}_{\text{inv}}=\bigcap_{j,k\in\mathbb{N}}\bigcup_{l\in\mathbb{N}}\left\{T\in\mathcal{T}:\,\|f_{j}-Tf_{l}\|<\frac{1}{k}\right\}=:\bigcap_{j,k\in\mathbb{N}}\bigcup_{l\in\mathbb{N}}M_{j,k,l}.$
It is clear that every set $M_{j,k,l}$ is open for the strong operator topology, finishing the proof.
4. Second Proof of $G_{\delta}$ Property
We also provide an alternative proof of the $G_{\delta}$ property via the points of continuity method.
The following is an ergodic theoretic analogue of Grivaux, Matheron, Menet [10, Prop. 2.11]. Here and later we denote by SOT and $\texttt{SOT}^{*}$ the strong and the strong operator topology on the Koopman operators, respectively.
Lemma 4.1.
The set $\mathcal{T}_{\text{inv}}$ coincides with the set of continuity points of the identity map $\text{id}:(\mathcal{T},\texttt{SOT})\to(\mathcal{T},\texttt{SOT}^{*})$.
Proof.
Let $T\in\mathcal{T}_{\text{inv}}$ and $(T_{n})\subset\mathcal{T}$ converging to $T$ in the strong operator topology. Then it converges also in the weak operator topology implying $T_{n}^{*}\to T^{*}$ weakly. Since $T$ is by assumption invertible, $T^{*}$ is an isometry as well and hence [^5] Recall that weak (operator) convergence of contractions on a Hilbert space implies strong convergence as soon as the limit operator is an isometry. $T_{n}^{*}\to T^{*}$ strongly. Thus $T$ is a point of continuity of $\text{id}:(\mathcal{T},\texttt{SOT})\to(\mathcal{T},\texttt{SOT}^{*})$.
Let now $T\in\mathcal{T}$ be such a point of continuity and let $f\in L^{2}(X,\mu)$. By Lemma 2.1 there exists a sequence $(T_{n})\subset\mathcal{T}_{\text{inv}}$ converging to $T$ weakly and hence strongly. By the point of continuity assumption we have
$\|f\|=\lim_{n\to\infty}\|T_{n}^{*}f\|=\|T^{*}f\|,$
and in particular $T^{*}$ is injective. Equivalently, $T$ has dense range, so [^6] Recall that isometries always have closed range. is surjective and hence invertible. This shows $T\in\mathcal{T}_{\text{inv}}$. ∎
The idea of the following proof in the operator theoretic setting, being a simplification of the one by Grivaux, Matheron, Menet [9, beginning of Sect. 3], was communicated to the author by Valentin Gillet.
Lemma 4.2.
The space $(\mathcal{T},\texttt{SOT}^{*})$ is separable.
Proof.
Let $(f_{n})_{n=1}^{\infty}$ be dense in $L^{2}(X,\mu)$ and consider the set
$A:=\{((Tf_{k})_{k=1}^{\infty},(T^{*}f_{k})_{k=1}^{\infty}):\,T\in\mathcal{T}\}\subset(L^{2}(X,\mu))^{\mathbb{N}}\times(L^{2}(X,\mu))^{\mathbb{N}}=:Y.$
Observe first that $Y$ is a Polish space, where $(L^{2}(X,\mu))^{\mathbb{N}}$ is endowed with the product topology induced by the norm topology on $L^{2}(X,\mu)$. Since for metric spaces separability is equivalent to second countability, $A$ is separable as well.
Thus let $(a_{n})$ be dense in $A$ and denote by $(T_{n})$ the corresponding countable subset of $\mathcal{T}$. We show that $(T_{n})$ is dense in $(\mathcal{T},\texttt{SOT}^{*})$. Let $T\in\mathcal{T}$ and let $(n_{l})$ be a subsequence of $\mathbb{N}$ such that $(a_{n_{l}})$ approximates $((Tf_{k})_{k=1}^{\infty},(T^{*}f_{k})_{k=1}^{\infty})$ in $Y$. This means that $(T_{n_{l}})$ approximates $T$ and $(T^{*}_{n_{l}})$ approximates $T^{*}$ in SOT, i.e., $(T_{n_{l}})$ approximates $T$ in $\texttt{SOT}^{*}$. The proof is complete. ∎
Proof of Theorem 1.1.
By a classical result of Baire, see Kechris [14, Thm. 24.14], the set of continuity points of a Baire class $1$ map [^7] A map $f:X\to Y$ between topological spaces $X$ and $Y$ is called of Baire class $1$ (or Borel $1$) if the preimage of every open subset of $Y$ is $F_{\sigma}$ in $X$. $f:X\to Y$ form a dense $G_{\delta}$ subset of $X$ whenever $X,Y$ are metrizable and $Y$ is separable. By Lemmata 4.1 and 4.2 it just remains to check the Baire class $1$ property of the identity map $\mathrm{id}:(\mathcal{T},\texttt{SOT})\to(\mathcal{T},\texttt{SOT}^{*})$. This can be done as in [10, Lemma 2.9] and we repeat here the simple argument for completeness.
Every $\texttt{SOT}^{*}$-open subset of $\mathcal{T}$ can be written as a countable union of finite intersections of sets either of the form
$U_{S,f,\varepsilon}:=\left\{T\in\mathcal{T}:\,\|(T-S)f\|<\varepsilon\right\}=\bigcup_{n=1}^{\infty}\left\{T\in\mathcal{T}:\,\|(T-S)f\|\leq\varepsilon-\frac{1}{n}\right\}$
or of the form
$U^{*}_{S,f,\varepsilon}:=\{T\in\mathcal{T}:\,\|(T^{*}-S^{*})f\|<\varepsilon\}=\bigcup_{n=1}^{\infty}\left\{T\in\mathcal{T}:\,\|(T^{*}-S^{*})f\|\leq\varepsilon-\frac{1}{n}\right\}$
for $S\in\mathcal{T}$, $f\in L^{2}(X,\mu)$ and $\varepsilon>0$. Since $\|(T^{*}-S^{*})f\|\leq\varepsilon-\frac{1}{n}$ can be written as
$|\langle(T-S)g,f\rangle|=|\langle g,(T^{*}-S^{*})f\rangle|\leq\varepsilon-\frac{1}{n}\qquad\forall g\in L^{2}(X,\mu)\text{ with }\|g\|=1,$
the SOT-$F_{\sigma}$ property of $\texttt{SOT}^{*}$-open sets is clear. Thus $\mathrm{id}:(\mathcal{T},\texttt{SOT})\to(\mathcal{T},\texttt{SOT}^{*})$ is of Baire class $1$. ∎